# UniST: A Prompt-Empowered Universal Model for Urban Spatio-Temporal Prediction

### v3

@@ 4.1. Base Model
The base model is a Transformer-based encoder-decoder architecture. Through spatio-temporal tokenizers, the base model can deal with diverse spatio-temporal data in a unified sequential format.
4.1.1. Spatio-temporal Tokenizers
The conventional Transformer architecture is designed for processing 1D sequential data. However, spatio-temporal data possesses a 4D structure. To accommodate this, we first split the data into channel-independent instances, which are 3D tensors. Then, we utilize spatio-temporal tokenizers to transform the 3D tensor, denoted as $X\in\mathbb{R}^{L\times H\times W}$, into multiple smaller 3D tensors. If the original shape is $L\times H\times W$, and the patch size is $(l,h,w)$, the resulting sequence is given by $E_{x}\in\mathbb{R}^{L^{\prime}\times H^{\prime}\times W^{\prime}},L^{\prime}=\frac{L}{l},H^{\prime}=\frac{H}{h},W^{\prime}=\frac{W}{w}$.
This transformation involves a 3D convolutional layer with a kernel size and stride both set to $(l,h,w)$. The process can be expressed as $E_{x}=\text{Conv}_{3d}(X))$, where $E_{x}$ represents the converted 1D sequential data. The sequence length of $E_{x}$ is $L^{\prime}\times H^{\prime}\times W^{\prime}$.
4.1.2. Positional Encoding.
As the original Transformer architecture does not consider the order of the sequence, we follow the common practice (Devlin et al., 2018) that incorporate positional encoding. To enhance generalization, we choose sine and cosine functions rather than learnable parameters for positional encoding. This encoding is separately applied to the spatial and temporal dimensions.
4.1.3. Encoder-decoder Structure
The base model adopts an encoder-decoder structure, following the methodology in Masked Autoencoder (MAE) (He et al., 2022). Input patches undergo masking at a certain ratio, and subsequently, the encoder receives the unmasked patches. Then, the decoder receives the encoder output and the masked patches and performs the reconstruction. However, our primary focus lies in capturing spatio-temporal dependencies encompassing both high-level and low-level relationships. For instance, we aim to predict precise values at specific grids in both time and space. Thus, we employ a decoder of comparable size to the encoder, different from the lightweight decoder used in MAE. Unlike the MAE decoder, exclusively utilized in the pre-training phase for image reconstruction, our decoder plays a pivotal role in both the pre-training and fine-tuning stages. The encoder and the decoder can be formulated as follows:
(1) $E_{enc}=\text{Encoder}(E_{x}),\ Y_{dec}=\text{Decoder}(E_{enc},E_{mask}),$
@@ 4.1.3. Encoder-decoder Structure
4.2. Spatio-temporal Self-supervised Pretraining
In pretrained language models, the self-supervised learning task is either masking-reconstruction (Devlin et al., 2018) or autoregressive prediction (Brown et al., 2020). Similarly, in vision models, visual patches are randomly masked and the pretraining objective is to reconstruct the masked pixels. To further augment the model’s capacity to capture intricate spatio-temporal relationships and intertwined dynamics, we introduce four distinct masking strategies during the pretraining phase, which are shown in the left box in the stage 1 of Figure 2. Suppose the masking percentage is $r$, we explain these strategies as follows:
@@ 4.2. Spatio-temporal Self-supervised Pretraining
    (2) $M\sim\textbf{U}[0,1],\ E_{x}=E_{x}[M<1-r],\ M\in\mathbb{R}^{L^{\prime}\times H^{\prime}\times W^{\prime}}.$
@@
    (3) $M\sim\textbf{U}[0,1],\ E_{x}=E_{x}[\ :\ ,M<1-r],\ M\in\mathbb{R}^{H^{\prime}\times W^{\prime}}.$
@@
    (4) $M\sim\textbf{Uniform}(1,2),\ E_{x}=E_{x}[\ :\ ,\frac{M-1}{2}H^{\prime}:\frac{M}{2}H^{\prime}].$
@@
    (5) $M=\text{Concat}([\textbf{1}_{(1-r)L^{\prime}\times H^{\prime}\times W^{\prime}},\textbf{0}_{rL^{\prime}\times H^{\prime}\times W^{\prime}}]),\ E_{x}=E_{x}[M=1].$
@@ 4.3. Spatio-temporal Knowledge-Guided Prompt
Prompt tuning plays a critical role in enhancing UniST’s generalization ability. Before delving into the details of our prompt design, it is essential to discuss why pre-trained models can be applied to unseen tasks and domains.
@@
In urban areas, the distributions of features and labels differ across cities and domains, denoted as $X_{A}\neq X_{B},Y_{A}\neq Y_{B}$, where $X$ and $Y$ denote features and labels, while $A$ and $B$ represent different cities or domains. Taken $A$ and $B$ as a simple example, generalization involves leveraging knowledge acquired from the $A$ dataset and adapt it to the $B$ dataset. The key point lies in identifying and aligning “related” patterns between $A$ and $B$ datasets. While finding similar patterns for an entire dataset may be challenging, we claim that identifying and aligning fine-grained patterns is feasible. Specifically, we provide some assumptions that applies to prompt-empowered spatio-temporal generalization, which are expressed as follows:
@@
(7) $P_{i}^{*}\neq P_{j}^{*}\quad\text{if}\quad D(x_{i},x_{j})>\epsilon,$
@@ 4.3.2. Spatio-Temporal Domain Knowledge
(8) $E_{sc}=\text{Conv}_{2D}[3](X_{s}),E_{sh}=\{\text{Conv}_{2D}[2^{i}+1](X_{s})\},i\in\{2,3,4\},$
(9) $E_{tc}=\text{Attention}_{t}(X_{c}),E_{tp}=\text{Attention}_{p}(X_{p}).$
It is essential to emphasize that the learning of $E_{sc},E_{sh},E_{tc}$, and $E_{tp}$ is not restricted by our practice. Practitioners have the flexibility to employ more complex designs to capture richer spatio-temporal properties. For example, Fourier-based approaches (Wu et al., 2022; Liu et al., 2023c) can be utilized to capture periodic patterns.
Figure 3. Illustration of the prompt generation process.
@@ 4.3.3. Spatio-temporal Prompt Learner
Given the representations of properties derived from spatio-temporal domain knowledge, the pivotal question is how to generate prompts—how does spatio-temporal knowledge guide prompt generation? Here we utilize prompt tuning techniques. While prompt tuning in computer vision (Jia et al., 2022) often train fixed prompts for specific tasks such as segmentation, detection, and classification. Due to the high-dimensional and complex nature of spatio-temporal patterns, training a fixed prompt for each case becomes impractical.
To tackle this issue, we draw inspirations from memory networks (Sukhbaatar et al., 2015) and propose a novel approach that learns a spatial memory pool and a temporal memory pool. In the prompt tuning process, these memory pools are optimized to store valuable information about spatio-temporal domain knowledge. As shown in Figure 3, the spatial and memory pools are defined as follows:
(10) $KM_{s}=\{(k_{s,0},m_{s,0}),(k_{s,1},m_{s,1}),...,(k_{s,N-1},m_{s,N-1})\},$
@@
Reduction 1.9% 0.42% 22.0% 15.7% 13.2% 20.6% 10.9% 21.6% 1.1% 7.1% 4.1% 13.3% 10.3% 13.6%
5. Results
@@ 5.1. Experimental Setup
To evaluate the performance of UniST, we conducted extensive experiments on more than 20 spatio-temporal datasets.
Datasets. The datasets we used cover multiple cities, spanning various domains such as crowd flow, dynamic population, traffic speed, cellular network usage, taxi trips, and bike demand. Appendix A Table 3 and Table 4 provide a summary of the datasets we used. These spatio-temporal datasets originate from distinct domains and cities, and have variations in the number of variables, sampling frequency, spatial scale, temporal duration, and data size.
@@
  • Deep urban prediction approaches. We consider state-of-the-art urban ST prediction models, including STResNet (Zhang et al., 2017), ACFM (Liu et al., 2018), MC-STL (Zhang et al., 2023b), STGSP (Zhao et al., 2022), STNorm (Deng et al., 2021), and STID (Shao et al., 2022).
  • Video prediction approaches. We compare with competitive video prediction models from the popular benchmark, including PredRNN (Wang et al., 2017), MAU (Chang et al., 2021), MIM (Wang et al., 2019), SimVP (Gao et al., 2022a), and TAU (Tan et al., 2023).
  • Multivariate time series forecasting approaches. We consider state-of-the-art multivariate time series forecasting models, including PatchTST (Nie et al., 2022) and iTransformer (Liu et al., 2023b). We also train PatchTST for all datasets, denoted as PatchTST(one-for-all).
  • Urban transfer learning approaches. We consider MAML (Finn et al., 2017) and MetaST (Yao et al., 2019) for evaluations of the generalization ability.
Metrics. To measure the prediction performance, we employed commonly used regression metrics, including Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE). For more detailed information on the datasets, baselines, and metrics used in our experiments, please refer to Appendix AC, and  D.
@@ 5.2. Short-Term Prediction
In short-term prediction, both the input step and prediction horizon are set as 6 following (Jin et al., 2023c; Nie et al., 2022). For baselines, we train a dedicated model for each dataset, while UniST is evaluated directly across all datasets.
Table 2. Performance comparison of long-term prediction on three datasets in terms of MAE and RMSE.
@@
Reduction 1.7% 8.2% 13.4% 14.0% 8.4% 14.9%
@@ Results.
Table 1 presents the short-term prediction results, with a selection of datasets due to space constraints. The complete results can be found in Table 6 and Table 7 in Appendix E. As we can observe from Table 1, UniST consistently outperforms all baselines across all datasets. Compared with the best baseline of each dataset, it showcases a notable average improvement of 11.3%. Notably, time series approaches such as PatchTST and iTransformer exhibit inferior performance compared to spatio-temporal methods. This underscores the importance of incorporating spatial dependency as prior knowledge for spatio-temporal prediction tasks. Another observation is that PatchTST(one-for-all) performs worth than PatchTST dedicated for each dataset, suggesting that the model struggles to directly adept to these distinct data distributions. Moreover, baseline approaches exhibit inconsistent performance across diverse datasets, indicating their instability across scenarios.
5.3. Long-Term Prediction
Setups.
Here we extend the input step and prediction horizon to 64 following (Jin et al., 2023c; Nie et al., 2022). This configuration accommodates prolonged temporal dependencies, allowing us to gauge the model’s proficiency in capturing extended patterns over time. Similar to the short-term prediction, UniST is directly evaluated across all datasets, while specific models are individually trained for each baseline on respective datasets.
Results.
Table 2 shows the long-term prediction results. Even with a more extended prediction horizon, UniST still consistently outperforms all baseline approaches across all datasets. Compared with the best baseline of each dataset, it yields an average improvement of 10.1%. This underscores UniST’s adept understanding of temporal patterns and its robust ability to generalize to extended timeframes. Table 8 in Appendix E illustrates the complete results.
@@ 5.4. Few-Shot Prediction
The hallmark of large foundation models lies in their exceptional generalization ability. The few-shot and zero-shot evaluations are commonly employed to characterize the ultimate tasks for universal time series forecasting (Zhou et al., 2023c; Jin et al., 2023c; Chang et al., 2023; Yuan et al., 2024). Likewise, the few-shot and zero-shot prediction capability is crucial for a universal spatio-temporal model. In this section, we assess the few-shot learning performance of UniST. Each dataset is partitioned into three segments: training data, validation data, and test data. In few-shot learning scenarios, when confronted with an unseen dataset during the training process, we utilized a restricted amount of training data, specifically, 1%, 5%, 10% of the training data. We choose some baselines with relatively good performance for the few-shot setting evaluation, We also compare with meta-learning baselines, i.e., MAML and MetaST, and pretraining and finetuning-based time series method, i.e., PatchTST.
@@ Results.
Appendix Table 9 to 11 illustrates the overall few-shot results. Due to the space limit, Figure 4 only illustrates the 1% few-shot learning results on two datasets. UniST still outperforms all baselines in few-shot learning scenarios, UniST achieves a larger relative improvement over baselines compared to long-term and short-term predictions. The transferability can be attributed to successful knowledge transfer in our spatio-temporal prompt.

### v4

@@ 3.3. Base Model
Our base model is a Transformer-based encoder-decoder architecture. Through spatio-temporal patching, it can handle diverse spatio-temporal data in a unified sequential format.
Spatio-Temporal Patching. The conventional Transformer architecture is designed for processing 1D sequential data. However, spatio-temporal data possesses a 4D structure. To accommodate this, we first split the data into channel-independent instances, which are 3D tensors. Then, we utilize spatio-temporal patching to transform the 3D tensor, denoted as $X\in\mathbb{R}^{L\times H\times W}$, into multiple smaller 3D tensors. If the original shape is $L\times H\times W$, and the patch size is $(l,h,w)$, the resulting sequence is given by $E_{x}\in\mathbb{R}^{L^{\prime}\times H^{\prime}\times W^{\prime}},L^{\prime}=\frac{L}{l},H^{\prime}=\frac{H}{h},W^{\prime}=\frac{W}{w}$.
This transformation involves a 3D convolutional layer with a kernel size and stride both set to $(l,h,w)$. The process can be expressed as $E_{x}=\text{Conv}_{3d}(X)$, where $E_{x}$ represents the converted 1D sequential data. The sequence length of $E_{x}$ is $L^{\prime}\times H^{\prime}\times W^{\prime}$.
Positional Encoding. As the original Transformer architecture does not consider the order of the sequence, we follow the common practice that incorporate positional encoding (Devlin et al., 2018). To enhance generalization, we choose sine and cosine functions rather than learnable parameters for positional encoding. This encoding is separately applied to the spatial and temporal dimensions.
Encoder-Decoder Structure. The base model utilizes an encoder-decoder framework inspired by Masked Autoencoder (MAE) (He et al., 2022). It processes input patches with a certain masking ratio, where the encoder takes the unmasked patches and the decoder reconstructs the image using the encoder’s output and the masked patches. Our focus is on capturing comprehensive spatio-temporal dependencies, including both high-level and low-level relationships, with the goal of accurately predicting values at specific time and space coordinates. Unlike MAE, which uses a lightweight decoder for pre-training, our model employs a full-sized decoder that plays a crucial role in both pre-training and fine-tuning. It can be formulated as:
$E_{enc}=\texttt{Encoder}(E_{x}),\ Y_{dec}=\texttt{Decoder}(E_{enc},E_{mask}),$
@@
3.4. Spatio-Temporal Self-Supervised Pre-train
In pretrained language models, the self-supervised learning task is either masking-reconstruction (Devlin et al., 2018) or autoregressive prediction (Brown et al., 2020). Similarly, in vision models, visual patches are randomly masked and the pre-training objective is to reconstruct the masked pixels. To further augment the model’s capacity to capture intricate spatio-temporal relationships and intertwined dynamics, we introduce four distinct masking strategies during the pre-training phase, which are shown in the left box in the stage 1 of Figure 2. Suppose the masking percentage is $r$, we explain these strategies as follows:
@@ 3.4. Spatio-Temporal Self-Supervised Pre-train
    $M\sim\textbf{U}[0,1],\ E_{x}=E_{x}[M<1-r],\ M\in\mathbb{R}^{L^{\prime}\times H^{\prime}\times W^{\prime}}.$
@@
    $M\sim\textbf{U}[0,1],\ E_{x}=E_{x}[\ :\ ,M<1-r],\ M\in\mathbb{R}^{H^{\prime}\times W^{\prime}}.$
@@
    $M\sim\textbf{{Uniform}}(1,2),\ E_{x}=E_{x}[\ :\ ,\frac{M-1}{2}H^{\prime}:\frac{M}{2}H^{\prime},\frac{M-1}{2}W^{\prime}:\frac{M}{2}W^{\prime}].$
@@
    $M=\texttt{Concat}([\textbf{1}_{(1-r)L^{\prime}\times H^{\prime}\times W^{\prime}},\textbf{0}_{rL^{\prime}\times H^{\prime}\times W^{\prime}}]),\ E_{x}=E_{x}[M=1].$
@@ 3.5. Spatio-Temporal Knowledge-Guided Prompt
Prompt learning plays a critical role in enhancing UniST’s generalization ability. Before delving into the details of our prompt design, it is essential to discuss why pre-trained models can be applied to unseen scenarios.
@@
In urban prediction tasks, the distributions of features and labels differ across domains and cities, denoted as $X_{A}\neq X_{B},Y_{A}\neq Y_{B}$, where $X$ and $Y$ denote features and labels, while $A$ and $B$ represent different cities or domains. Taken $A$ and $B$ as a simple example, generalization involves leveraging knowledge acquired from the $A$ dataset and adapt it to the $B$ dataset. The key point lies in identifying and aligning “related” patterns between $A$ and $B$ datasets. While finding similar patterns for an entire dataset may be challenging, we claim that identifying and aligning fine-grained patterns is feasible. Specifically, we provide some assumptions that applies to prompt-empowered spatio-temporal generalization, which are expressed as follows:
@@
$P_{i}^{*}\neq P_{j}^{*}\quad\text{if}\quad D(x_{i},x_{j})>\epsilon,$
@@ 3.5.2. Spatio-Temporal Domain Knowledge
$E_{sc}=\texttt{Conv}_{2D}[3](X_{s}),E_{sh}=\{\texttt{Conv}_{2D}[2^{i}+1](X_{s})\},i\in\{2,3,4\},$
$E_{tc}=\texttt{Attention}(X_{c}),E_{tp}=\texttt{Attention}(X_{p}).$
It is essential to emphasize that the learning of $E_{sc},E_{sh},E_{tc}$, and $E_{tp}$ is not restricted by our practice. Practitioners have the flexibility to employ more complex designs to capture richer spatio-temporal properties. For example, Fourier-based approaches (Wu et al., 2022; Liu et al., 2023b) can be utilized to capture periodic patterns.
@@ 3.5.3. Spatio-Temporal Prompt Learner
Given the representations of properties derived from spatio-temporal domain knowledge, the pivotal question is how to generate prompts—how does spatio-temporal knowledge guide prompt generation? Here we utilize prompt learning techniques. While prompt learning in computer vision (Jia et al., 2022) often train fixed prompts for specific tasks such as segmentation, detection, and classification. Due to the high-dimensional and complex nature of spatio-temporal patterns, training a fixed prompt for each case becomes impractical.
To tackle this issue, we draw inspirations from memory networks (Sukhbaatar et al., 2015) and propose a novel approach that learns a spatial memory pool and a temporal memory pool. In the prompt learning process, these memory pools are optimized to store valuable information about spatio-temporal domain knowledge. As shown in Figure 3, the spatial and memory pools are defined as follows:
$KM_{s}=\{(k_{s,0},m_{s,0}),(k_{s,1},m_{s,1}),...,(k_{s,N-1},m_{s,N-1})\},$
@@
$\alpha_{sc}=[k_{s,0};k_{s,1};...,k_{s,N-1}]E_{sc}^{T},\ P_{sc}=\sum_{i}\alpha_{sc,i}m_{s,i},$
4. Performance Evaluations
@@ 4.1. Experimental Setup
To evaluate the performance of UniST, we conducted extensive experiments on more than 20 spatio-temporal datasets. Due to the page limit, we select a few representative results on below, and a full benchmark can be found in the arXiv version [^1] https://arxiv.org/abs/2402.11838.
Datasets. The datasets we used cover multiple cities, spanning various domains such as crowd flow, dynamic population, traffic speed, cellular network usage, taxi trips, and bike demand. Appendix Table 4 and Table 5 (arXiv) provide a summary of the datasets we used. These spatio-temporal datasets originate from distinct domains and cities, and have variations in the number of variables, sampling frequency, spatial scale, temporal duration, and data size.
@@
  • Deep urban prediction approaches. We consider state-of-the-art urban ST prediction models, including STResNet (Zhang et al., 2017), ACFM (Liu et al., 2018), MC-STL (Zhang et al., 2023a), STGSP (Zhao et al., 2022), STNorm (Deng et al., 2021), STID (Shao et al., 2022b), and PromptST (Zhang et al., 2023c).
  • Video prediction approaches. We compare with competitive video prediction models from the popular benchmark, including PredRNN (Wang et al., 2017), MAU (Chang et al., 2021), MIM (Wang et al., 2019), SimVP (Gao et al., 2022), and TAU (Tan et al., 2023).
  • Multivariate time series forecasting approaches. We consider state-of-the-art multivariate time series forecasting models, including PatchTST (Nie et al., 2022) and iTransformer (Liu et al., 2023a). For a fair comparison, we also train PatchTST for all datasets, denoted as PatchTST(one-for-all).
  • Meta learning approaches. To evaluate the generalization capability, we consider meta-learning approaches including MAML (Finn et al., 2017) and MetaST (Yao et al., 2019).
Metrics. We employed commonly used regression metrics, including Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE). For more detailed information of the datasets, baselines, and metrics, please refer to Appendix AB, and Appendix D (arXiv).
@@ 4.2. Short-Term Prediction
Following previous practices (Jin et al., 2023a; Nie et al., 2022), both the input step and prediction horizon are set as 6, i.e., $6\rightarrow 6$. . For baselines, we train a dedicated model for each dataset, while UniST is evaluated across all datasets.
Results.
Table 2 presents the short-term prediction results, with a selection of datasets due to space constraints. The complete results can be found in Table 11 and Table 12 in Appendix E (arXiv). As we can observe from Table 2, UniST consistently outperforms all baselines across all datasets. Compared with the best baseline of each dataset, it showcases a notable average improvement. Notably, time series approaches such as PatchTST and iTransformer exhibit inferior performance compared to spatio-temporal methods. This underscores the importance of incorporating spatial dependency as prior knowledge for spatio-temporal prediction tasks. Another observation is that PatchTST(one-for-all) performs worse than PatchTST dedicated for each dataset, suggesting that the model struggles to directly adept to these distinct data distributions. Moreover, baseline approaches exhibit inconsistent performance across diverse datasets, indicating their instability across scenarios. The consistent superior performance of UniST across all scenarios underscores the significant potential and benefits of a one-for-all model. Moreover, it demonstrates UniST’s capability to orchestrate diverse data, where different datasets can benefit each other.
4.3. Long-Term Prediction
Table 3. Performance comparison of long-term prediction on three datasets in terms of MAE and RMSE. We use the average prediction errors over all prediction steps. Bold denotes the best results and underline denotes the second-best results.
@@ 4.3. Long-Term Prediction
Setups.
Here we extend the input step and prediction horizon to 64 following (Jin et al., 2023a; Nie et al., 2022). This configuration accommodates prolonged temporal dependencies, allowing us to gauge the model’s proficiency in capturing extended patterns over time. Similar to the short-term prediction, UniST is directly evaluated across all datasets, while specific models are individually trained for each baseline on respective datasets.
@@ Results.
Table 3 shows the long-term prediction results. Even with a more extended prediction horizon, UniST still consistently outperforms all baseline approaches across all datasets. Compared with the best baseline of each dataset, it yields an average improvement of 10.1%. This highlights UniST’s capability to comprehend temporal patterns effectively and its robustness in generalizing across extended durations. Table 13 in Appendix E (arXiv) illustrates the complete results.
@@ 4.4. Few-Shot Prediction
The hallmark of large foundation models lies in their exceptional generalization ability. The few-shot and zero-shot evaluations are commonly employed to characterize the ultimate tasks for universal time series forecasting (Zhou et al., 2023b; Jin et al., 2023a; Chang et al., 2023). Likewise, the few-shot and zero-shot prediction capability is crucial for a universal spatio-temporal model. In this section, we assess the few-shot learning performance of UniST. Each dataset is partitioned into three segments: training data, validation data, and test data. In few-shot learning scenarios, when confronted with an unseen dataset during the training process, we utilized a restricted amount of training data, specifically, 1%, 5%, 10% of the training data. We choose some baselines with relatively good performance for the few-shot setting evaluation, We also compare with meta-learning baselines, i.e., MAML and MetaST, and pretraining and finetuning-based time series method, i.e., PatchTST.
@@ Results.
Appendix Table 14 to 16 (arXiv) illustrates the overall few-shot results. Due to the space limit, Figure 4 only illustrates the 1% few-shot learning results on two datasets. In these cases, UniST still outperforms all baselines, it achieves a larger relative improvement over baselines compared to long-term and short-term predictions. The transferability can be attributed to successful knowledge transfer in our spatio-temporal prompt.
@@ 4.5. Zero-Shot Prediction
Results.
Figure 4 also compare the performance of UniST (zero-shot) and baselines (few-shot). As observed, UniST achieves remarkable zero-shot performance, even surpassing many baselines trained with training data that are highlighted by red dashed lines. We attribute these surprising results to the powerful spatio-temporal transfer capability. It suggests that for a completely new scenario, even when the displayed overall patterns are dissimilar to the data encountered during the training process, UniST can extract fine-grained similar patterns from our defined spatial and temporal properties. The few-shot and zero-shot results demonstrate the powerful generalization capability of UniST.
5. Study and Aalysis on UniST